### v2

# Brain-Inspired Stepwise Patch Merging for Vision Transformers

Abstract
The hierarchical architecture has become a mainstream design paradigm for Vision Transformers (ViTs), with Patch Merging serving as the pivotal component that transforms a columnar architecture into a hierarchical one. Drawing inspiration from the brain’s ability to integrate global and local information for comprehensive visual understanding, we propose Stepwise Patch Merging (SPM), which enhances the subsequent attention mechanism’s ability to ’see’ better. SPM consists of Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE) striking a proper balance between long-range dependency modeling and local feature enhancement. Extensive experiments conducted on benchmark datasets, including ImageNet-1K, COCO, and ADE20K, demonstrate that SPM significantly improves the performance of various models, particularly in dense prediction tasks such as object detection and semantic segmentation. Meanwhile, experiments show that combining SPM with different backbones can further improve performance. The code has been released at https://github.com/Yonghao-Yu/StepwisePatchMerging.
1 Introduction
Transformers have demonstrated remarkable advancements in natural language processing (NLP) Vaswani et al. (2017); Devlin et al. (2018), and their application has recently extended significantly into the computer vision (CV) domain Dosovitskiy et al. (2020). To enhance their adaptability to downstream tasks, hierarchical vision transformers (HVTs) Wang et al. (2021); Liu et al. (2021) have been developed. These architectures draw inspiration from the pyramid structure utilized in convolutional neural networks (CNNs) Krizhevsky et al. (2012); He et al. (2016). In HVTs, transformer blocks are segmented into multiple stages, resulting in a progressive reduction of feature map sizes and an increase in the number of channels as the network depth increases.
Figure 1: Overview of the proposed Stepwise Patch Merging (SPM) framework built upon the HVTs. The SPM framework comprises two sequential modules: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE).
HVTs commonly utilize either standard convolutional layers or linear projection layers to amalgamate adjacent tokens Wang et al. (2021); Liu et al. (2021); Ren et al. (2022), with the objective of generating hierarchical feature maps. Nevertheless, fixed-grid methods can limit the representational capacity of vision transformers in modeling geometric transformations, as not every pixel contributes equally to an output unit Luo et al. (2016). To overcome this limitation, adaptive methods have been proposed to derive more informative downsampled tokens for subsequent processing. For example, LIT Pan et al. (2022), drawing inspiration from deformable convolutions Dai et al. (2017), learns a grid of offsets to adaptively adjust spatial sampling locations for merging neighboring patches from a sub-window in a feature map. Similarly, TCformer Zeng et al. (2022) employs a variant of the k-nearest neighbor-based density peaks clustering algorithm (DPC-KNN) Du et al. (2016) to aggregate redundant patches, generating more patches on the target object to capture additional information. HAFA Chen et al. (2023) integrates the methodologies of LIT and TCformer, predicting offsets to adjust the sampling center of patches in the shallow layers, while employing clustering in the deeper layers to group patches with similar semantics in the feature space. However, these methods face several common challenges. They often exhibit limited capacity for modeling long-distance relationships and suffer from a loss of spatial information due to the clustering process. Additionally, the clustering algorithms used are typically not amenable to end-to-end training, leading to inefficiencies. The integration of multiple modules, as seen in HAFA, further complicates their generalizability across different applications.
In the brain, the primary visual cortex (V1), integral to initial visual processing, houses neurons with relatively small receptive fields that are crucial for detecting fine, localized visual features such as edges and orientations. As visual information propagates to higher cortical areas like V2, V3, and V4, neurons with increasingly larger receptive fields integrate these initial perceptions, facilitating the recognition of more complex patterns and broader contextual elements Hubel and Wiesel (1962). Additionally, the visual cortex benefits from a dynamic feedback system where higher-order areas like the inferotemporal cortex (IT) provide contextual modulation to lower areas. This top-down modulation is essential for refining the perception of local features within their broader environmental matrix, enhancing both the accuracy and relevance of visual processing Gilbert and Li (2013).
Inspired by the orchestrated activities across various cortical areas, we introduce Stepwise Patch Merging (SPM), as shown in Fig. 1, a novel approach designed to enhance the receptive field while preserving local details. SPM framework consists of two sequential stages: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE). In the MSA stage, spatial dimensions are preserved while channel dimensions are increased to a designated size. This process aggregates multi-scale information, enriching the semantic content to accommodate the increased capacity of the feature map. Subsequently, the GLE stage reduces the spatial dimensions of the feature map while maintaining the channel dimensions. Given that the input to GLE already contains rich semantic information, this stage emphasizes local information, optimizing it for downstream dense prediction tasks such as object detection and semantic segmentation. The distinct focus and reasonable division of labor between the MSA and GLE modules ensure that the SPM architecture serves as a flexible, drop-in replacement for existing HVTs.
In summary, our contributions are as follows:
  • We propose an innovative technique termed Stepwise Patch Merging (SPM), which serves as a plug-in replacement within HVTs, leading to substantial performance enhancements.
  • The SPM framework comprises two distinct modules: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE). MSA enriches feature representation by integrating multi-scale information, while GLE enhances the extraction of local details, achieving an optimal balance between long-range dependency modeling and local feature refinement.
  • Extensive experiments conducted on benchmark datasets, including ImageNet-1K, COCO, and ADE20K, demonstrate that SPM significantly boosts the performance of various models, particularly in downstream dense prediction tasks such as object detection and semantic segmentation.
2 Related Work
2.1 Vision Transformer
The Vision Transformer (ViT) Dosovitskiy et al. (2020) revolutionized visual tasks by introducing the transformer architecture to computer vision. ViT segments images into non-overlapping patches, projects these patches linearly into token sequences, and processes them using a transformer encoder. ViT models have demonstrated superior performance in image classification and other downstream tasks, surpassing CNNs He et al. (2016); Krizhevsky et al. (2012) when trained with large-scale pretraining datasets and advanced training methodologies. Motivated by the success of CNNs and the necessity to address dense prediction tasks, researchers have incorporated the feature pyramid structure within transformers. This innovation has led to the development and widespread adoption of HVTs Dong et al. (2022); Ren et al. (2022); Guo et al. (2022); Hou et al. (2022); Lin et al. (2023).
2.2 Hierarchical Feature Representation
Hierarchical feature representation plays a pivotal role in dense prediction tasks, prompting extensive research in this domain. Existing approaches can be broadly categorized into fixed-grid and dynamic feature-based methods. Fixed-grid methods, exemplified by works such as PVT Wang et al. (2021) and Swin Liu et al. (2021), merge patches within adjacent windows using 2D convolution. In contrast, dynamic methods, such as DynamicViT Rao et al. (2021) adaptively extract features by eliminating redundant patches and retaining essential ones, thereby forming hierarchical feature maps. EviT Liang et al. (2022) enhances this approach by selecting the top K tokens with the highest average values across all heads for the next stage, merging the remaining tokens. PS-ViT Yue et al. (2021) further refines the process by iteratively adjusting patch centers towards the object to enrich object information within the hierarchical feature maps. Token Merging Bolya et al. (2022) employs cosine similarity to progressively merge similar tokens, thereby increasing model throughput.
Fixed-grid methods are constrained by their singular and relatively small receptive fields, and excessively enlarging the grid size leads to increased computational overhead. Dynamic feature-based methods, while adaptive, may discard low-scoring tokens that contain valuable information and often lack end-to-end training capabilities. Our proposed SPM distinguishes itself from both fixed-grid and dynamic feature-based methods.
2.3 Global/Coarse and Local/Fine Feature Fusion
Fusing coarse- and fine-grained features is not novel; however, most prior work treats both feature types equally and performs simple fusion. For example, PLG-ViT Ebert et al. (2023) fuses global and local features using the Hadamard product. SPM uses a lightweight module (GTG) to generate global information that guides the generation of detailed features, an approach demonstrated by experiments to be highly effective.
3 Methodology
Inspired by the brain’s ability to integrate global and local information when processing visual scenes, we propose the Stepwise Patch Merging framework, as illustrated in Fig. 1. The framework comprises two primary components: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE), designed to address variations in feature map dimensions. The MSA module enhances feature diversity by increasing the number of channels and capturing long-range dependencies, akin to how the brain processes information at multiple scales to form a coherent perception. In contrast, the GLE module optimizes local feature extraction by introducing context-aware guide tokens within local windows, thereby refining and enhancing feature details. This synergistic design effectively combines the strengths of both global structure processing and local detail enhancement, making it particularly beneficial for downstream dense prediction tasks.
Figure 2: An illustration of Multi-Scale Aggregation (MSA).
3.1 Multi-Scale Aggregation
Our proposed Multi-Scale Aggregation (MSA) module draws inspiration from the brain’s remarkable ability to effectively model long-range dependencies when processing visual information. In the brain, the visual system achieves precise modeling of long-range dependencies through multi-level and multi-scale information processing. Neurons with small receptive fields process local features, and this information is progressively integrated over larger areas by neurons with larger receptive fields, capturing complex patterns and objects. Additionally, the brain’s extensive network of long-range neural connections allows for the exchange and integration of data from various parts of the visual field, facilitating a comprehensive understanding of the scene. Furthermore, neurons within the same level possess receptive fields of varying sizes, enabling the brain to simultaneously process local details and global features. This sophisticated mechanism of combining local and global information processing in the brain inspired the design of our MSA module, which aims to enhance feature diversity and capture long-range dependencies effectively.
Inspired by these mechanisms, the MSA module first divides the input channels $C$ into $N$ distinct heads, each undergoing depth-wise convolutions with varying receptive fields. This method not only reduces the parameter count and computational cost but also facilitates the extraction of multi-granularity information, akin to how different neurons in the brain handle information processing. Subsequently, the MSA module employs larger convolutional kernels to further expand the receptive field, thereby enhancing its capability to model long-range dependencies. Following this, Channel Shuffle Zhang et al. (2018) is used to interleave channels containing features of different scales, followed by a series of linear projections to fuse these multi-scale features. The number of linear projections is $\frac{C}{N}$, with each projection having unique parameters. Finally, the $N$ heads are concatenated, and a final linear projection adjusts the number of channels to the specified $C^{\prime}$.
By leveraging the brain’s mechanism for effective long-range dependency modeling, the MSA module better captures and integrates key features, significantly enhancing the model’s performance in complex visual recognition tasks.
Our proposed MSA can be formulated as follows:
$H_{n}=DWConv_{k_{n}\times k_{n}}(x_{n})$ (1)
$G^{c}=\mathbf{W}^{c}([H_{1}^{c};H_{2}^{c};...;H_{N}^{c}])$
$\bm{MSA}(X)=\mathbf{W}([G^{1};G^{2};...;G^{\frac{C}{N}}])$
where $X=[x_{1},x_{2},\ldots,x_{N}]$ represents the input $X$ split into multiple heads along the channel dimension, and $x_{n}\in\mathbb{R}^{B\times H\times W\times\frac{C}{N}}$ denotes the $n$-th head. The kernel size of the depth-wise convolution for the $n$-th head is denoted by $k_{n}\in{k_{1},k_{2},\ldots,k_{N}}$. Here, $H_{n}\in\mathbb{R}^{B\times H\times W\times\frac{C}{N}}$ represents the $n$-th head after being processed by the depth-wise convolution with $out_{c}hannels=in_{c}hannels$, and $H_{n}^{c}$ represents the $c$-th channel in the $n$-th head. $\mathbf{W}^{c}\in\mathbb{R}^{N\times N}$ is the weight matrix of the linear projection. $\mathbf{G}^{c}\in\mathbb{R}^{B\times H\times W\times N}$. Finally, $\mathbf{W}\in\mathbb{R}^{C\times C^{\prime}}$ is the weight matrix of the linear projection that adjusts the number of channels to the specified $C^{\prime}$.
3.2 Guided Local Enhancement
Inspired by the brain’s ability to enhance local features through context-aware processing, we developed the Guided Local Enhancement (GLE) module. In the brain, local feature enhancement is achieved by integrating information from both local and global contexts. Higher-level cortical areas provide contextual feedback that refines the processing of local features, ensuring that details are interpreted within the broader visual context. This hierarchical processing involves neurons that respond specifically to local stimuli but are influenced by surrounding contextual information, allowing for more nuanced and precise feature extraction.
Following this principle, the GLE module acts as a local feature enhancer utilizing context-aware guide tokens, as illustrated in Fig. 3. Specifically, we implement self-attention within a local window and introduce a $[Guide]$ token into the input space. This $[Guide]$ token undergoes the same self-attention operations as the patch tokens, thereby providing high-level semantic features to the local window during pooling. By mimicking the brain’s method of using contextual information to refine local feature extraction, the GLE module ensures that the extracted local features are both precise and contextually relevant.
Figure 3: An illustration of Guided Local Enhancement (GLE).
Formally, given an input $X\in\mathbb{R}^{C\times H\times W}$, the $[Guide]$ tokens are generated by a large-kernel depth-wise convolution, referred to as the Guide Token Generator (GTG), which can be described as follows:
$\bm{GTG}(X)=DWConv(GELU(BatchNorm(X))).$ (2)
We focus on the operations performed on a single pixel within the input feature map. We define a set of pixels within a local window centered at pixel $(i,j)$ as $\rho(i,j)$. For a fixed window size of $k\times k$, $\|\rho(i,j)\|=k^{2}$. In our setup, $k$ is equal to the stride of the GTG, both being 2, meaning #$\#Windows=\|GTG(X)\|=\frac{H\times W}{4}$. Tokens within a window containing a $[Guide]$ token can be represented by the sequence $\mathbf{z}$:
$\mathbf{z}=\left[S_{(i,j)\sim GTG(X)};S^{1}_{(i,j)\sim\rho(i,j)};\cdots;S^{k^{2}}_{(i,j)\sim\rho(i,j)}\right],$ (3)
then perform the standard self-attention operation on $\mathbf{z}$:
$\left[\mathbf{q},\mathbf{k},\mathbf{v}\right]$ $=\mathbf{z}\mathbf{U}_{qkv},$ (4)
$SA(\mathbf{z})$ $=softmax(\frac{\mathbf{q}\mathbf{k}^{\top}}{\sqrt{C}})\mathbf{v},$
where $\mathbf{U}_{qkv}\in\mathbb{R}^{C\times 3C}$, and we ignore the relative positional relationships between tokens within a window, so positional encoding is not used. Finally, we select the $[Guide]$ token as the output of the GLE:
$SA(\mathbf{z})=\left[A_{(i,j)\sim GTG(X)};A_{(i,j)\sim\rho(i,j)}\right],$ (5)
$\bm{GLE}(X_{(i,j)})=A_{(i,j)\sim GTG(X)}.$
It is worth noting that when the $[Guide]$ token is stripped of its semantic information, it degrades into the $[CLS]$ token in the vanilla vision transformer Dosovitskiy et al. (2020). Experiments show that using our $[Guide]$ tokens results in higher performance (see Tab. 6).
4 Experiments
4.1 Image Classification on ImageNet-1K
Setting.
We first evaluate the proposed SPM framework on the ImageNet-1K dataset Deng et al. (2009), which comprises 1.28 million training images and 50,000 validation images spanning 1,000 categories. To ensure a fair comparison, all models are trained on the training set and report the top-1 error rate on the validation set. For data augmentation, we apply a suite of techniques including random cropping, random horizontal flipping Szegedy et al. (2015), label smoothing regularization Szegedy et al. (2016), mixup Zhang et al. (2017), CutMix Yun et al. (2019), and random erasing Zhong et al. (2020). These augmentations are employed to enhance the robustness and generalization ability of the models. During training, we use the AdamW optimizer Loshchilov and Hutter (2017) with a momentum parameter of 0.9, a mini-batch size of 128, and a weight decay of $5\times 10^{-2}$. The initial learning rate is set to $1\times 10^{-3}$ and follows a cosine annealing schedule Loshchilov and Hutter (2016) to gradually reduce the learning rate. All models are trained from scratch for 300 epochs on eight NVIDIA A100 GPUs. For evaluation, we adopt the standard center crop strategy on the validation set, where a $224\times 224$ patch is extracted from each image to assess the classification accuracy.
Backbone #Params (M) Top-1 Acc. (%)
ResNet18 He et al. (2016) 11.7 68.5
DeiT-Tiny/16 Touvron et al. (2021) 5.7 72.2
PVT-Tiny Wang et al. (2021) 13.2 75.1
PVT-Tiny (HAFA) Chen et al. (2023) 14.6 77.5
PVT-Tiny (SPM) 14.0 79.5 (+4.4)
ResNet50 He et al. (2016) 25.6 78.5
ResNeXt50-32$\times$4d Xie et al. (2017) 25.0 79.5
DeiT-Small/16 Touvron et al. (2021) 22.1 79.9
HRNet-W32 Wang et al. (2020) 41.2 78.5
PVT-Small Wang et al. (2021) 24.5 79.8
PVT-Small (HAFA) Chen et al. (2023) 25.8 80.1
PVT-Small (SPM) 25.3 81.7 (+1.9)
ResNeXt101-64$\times$4d Xie et al. (2017) 83.5 81.5
ViT-Base/16 Dosovitskiy et al. (2020) 86.6 81.8
DeiT-Base/16 Touvron et al. (2021) 86.6 81.8
PVT-Medium Wang et al. (2021) 44.2 81.2
PVT-Medium (SPM) 45.0 81.9 (+0.7)
Table 1: Image classification performance on the ImageNet validation set. “#Params” refers to the number of parameters.
Backbone #Params (M) Mask R-CNN
$AP^{b}$ $AP_{50}^{b}$ $AP_{75}^{b}$ $AP_{s}^{b}$ $AP_{m}^{b}$ $AP_{l}^{b}$ $AP^{m}$ $AP_{50}^{m}$ $AP_{75}^{m}$
ResNet18 He et al. (2016) 31.2 34.0 54.0 36.7 - - - 31.2 51.0 32.7
PVT-Tiny Wang et al. (2021) 32.9 36.7 59.2 39.3 21.6 39.2 49.0 35.1 56.7 37.3
PVT-Tiny (HAFA) Chen et al. (2023) 34.5 39.8 62.6 43.3 23.3 42.7 53.3 37.1 59.4 39.3
PVT-Tiny (SPM) 33.7 40.8 (+4.1) 63.4 44.3 24.9 44.0 54.0 38.0 (+2.9) 60.4 40.6
ResNet50 He et al. (2016) 44.2 38.0 58.6 41.4 - - - 34.4 55.1 36.7
PVT-Small Wang et al. (2021) 44.1 40.4 62.9 43.8 22.9 43.0 55.4 37.8 60.1 40.3
PVT-Tiny (HAFA) Chen et al. (2023) 45.8 41.8 64.4 45.7 26.0 44.6 56.1 38.9 61.5 41.9
PVT-Small (SPM) 44.9 43.0 (+2.6) 65.4 46.7 25.5 46.2 57.6 39.6 (+1.8) 62.3 42.4
ResNet101 He et al. (2016) 63.2 40.4 61.1 44.2 - - - 36.4 57.7 38.8
ResNeXt101-32$\times$4d Xie et al. (2017) 62.8 41.9 62.5 45.9 - - - 37.5 59.4 40.2
PVT-Medium Wang et al. (2021) 63.9 42.0 64.4 45.6 - - - 39.0 61.6 42.1
PVT-Medium (SPM) 64.7 43.3 (+1.3) 64.9 47.6 25.8 46.4 58.3 39.4 (+0.4) 61.7 42.2
Table 2: Object detection and instance segmentation performance on COCO val2017. $AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
Result.
In Tab. 1, we observe that incorporating the SPM framework into the PVT results in significant improvements in classification accuracy, specifically by 4.4%, 1.9%, and 0.7% in the Tiny, Small, and Medium models, respectively. The final experimental results indicate that the accuracy of models of various sizes has been enhanced, with the most notable improvement observed in the Tiny model. Furthermore, with the integration of SPM, PVT-Small surpassed PVT-Medium by 0.5%.
4.2 Object Detection on COCO
Setting.
Object detection and instance segmentation experiments were conducted on the challenging COCO benchmark Lin et al. (2014). All models were trained on the training set comprising 118k images and evaluated on the validation set with 5k images. We validated the effectiveness of different backbones using Mask R-CNN He et al. (2017). Before training, the weights pre-trained on ImageNet-1K were used to initialize the backbone, and the newly added layers were initialized using the Xavier initialization method Glorot and Bengio (2010). Our models were trained with a batch size of 16 on 8 NVIDIA A100 GPUs and optimized using the AdamW optimizer Loshchilov and Hutter (2017) with an initial learning rate of $1\times 10^{-4}$.
Result.
As shown in Tab. 2, incorporating the SPM framework into the PVT resulted in significant improvements of 4.1%, 2.6%, and 1.3% in the Tiny, Small, and Medium models, respectively, for the object detection task. Models integrated with SPM show marked improvements in detecting medium-sized and large objects. This enhancement is attributed to the original patch merging’s relatively singular and small receptive fields, whereas the MSA module integrates features with diverse receptive fields, enabling the model to more accurately capture long-range dependencies. Moreover, there is a significant improvement in detecting small objects. Although larger models are typically better at modeling global relationships, the disruption of local information may hinder small objects from establishing complete semantic information, leading to missed detections. The GLE module addresses this by enhancing the perception of local discriminative information, resulting in consistent improvements in the detection performance of small objects across models of different sizes.
4.3 Semantic Segmentation on ADE20K
Backbone Semantic FPN
#Params (M) mIoU (%)
ResNet18 He et al. (2016) 15.5 32.9
PVT-Tiny Wang et al. (2021) 17.0 35.7
PVT-Tiny (HAFA) Chen et al. (2023) 18.7 40.1
PVT-Tiny (SPM) 17.8 41.5 (+5.8)
ResNet50 He et al. (2016) 28.5 36.7
PVT-Small Wang et al. (2021) 28.2 39.8
PVT-Small (HAFA) Chen et al. (2023) 29.9 43.8
PVT-Small (SPM) 29.0 45.9 (+6.1)
ResNet101 He et al. (2016) 47.5 38.8
ResNeXt101-32$\times$4d Xie et al. (2017) 47.1 39.7
PVT-Medium Wang et al. (2021) 48.0 41.6
PVT-Medium (SPM) 48.8 45.3 (+3.7)
Table 3: Semantic segmentation performance of different backbones on the ADE20K validation set.
Setting.
The ADE20K dataset Zhou et al. (2017) is a widely utilized benchmark for semantic segmentation, comprising 150 categories with 20,210 images for training, 2,000 images for validation, and 3,352 images for testing. All the methods compared were evaluated using the Semantic FPN framework Kirillov et al. (2019). The backbone network of our method was initialized with the pre-trained ImageNet-1k model, and the newly added layers were initialized using the Xavier initialization method. The initial learning rate was set to 0.0001, and the model was optimized using the AdamW optimizer. We trained our models for 40,000 iterations with a batch size of 16 on eight NVIDIA A100 GPUs. The learning rate followed a polynomial decay schedule with a power of 0.9. During training, images were randomly resized and cropped to $512\times 512$ pixels. For testing, images were rescaled to have a shorter side of 512 pixels.
Result.
As shown in Tab. 3, the integration of the SPM framework led to a significant enhancement in the semantic segmentation task. Specifically, the performance of the Tiny, Small, and Large models improved by 5.8%, 6.1%, and 3.7%, respectively. Interestingly, in both classification and detection tasks, the relative improvement brought by SPM compared to the base model gradually decreases as the model size increases. However, on the ADE20K dataset, the performance gain of PVT-Small exceeds that of PVT-Tiny, with improvements of 6.1% and 5.8%, respectively.
Figure 4: Visualization of the attention map includes the original images (the 1st column), the visualization of the attention map for each head after using SPM (the 2nd and 3rd columns), and without using SPM (the 4th and 5th columns). The red five-pointed star in the original images represents the position of the query.
4.4 Effectiveness on Other Backbones
To further validate the generalizability of the proposed SPM framework, we integrated SPM into various mainstream Transformer backbones and trained them on the ImageNet-1K dataset. We employed consistent training settings to ensure a fair comparison, and the top-1 accuracies are presented in Tab. 4. The results demonstrate that the performance improvements conferred by SPM are universal across different backbones, indicating its robust generalization capability. Specifically, our method significantly enhanced the performance of Swin-T by 1.1%, Shunted-T by 0.8%, and NAT-Mini by 0.4% on ImageNet-1K. These results underscore the effectiveness of SPM in boosting the performance of various Transformer architectures, highlighting its potential as a versatile enhancement technique in the field of computer vision.
4.5 Visualization
We compared the visualization results of the attention maps with and without using the SPM framework, as shown in Fig. 4. Specifically, we replaced the original patch merging block of PVT-Tiny with our SPM and visualized the first block of the second stage for two separate heads. For example, after employing SPM, the bird’s two wings and tail were successfully linked, whereas the vanilla PVT-Tiny failed to capture the distant tail. This demonstrates that SPM facilitates the network’s ability to establish long-range relationships at shallower layers, leading to significant improvements in classification performance.
As shown in Fig. 5 (a), MSA enhances fine-grained features and moderate global context benefits detail generation, while excessive global aggregation leads to performance degradation (see Tab. 7).
Additionally, we employed the Effective Receptive Field (ERF) method Luo et al. (2016) as a visualization tool to compare the changes in ERF before and after using SPM, as depicted in Fig. 5. It is readily observed that after integrating SPM, the size of the ERF not only increases significantly but also changes shape from a regular square to a radial decay pattern, which aligns more closely with biological vision.
Figure 5: (a) Visualizations before/after MSA and GLE with increasing GTG kernel size (3, 5 and 7); (b) Visualization comparison of the Effective Receptive Field of PVT-Tiny before (the 1st row) and after applying our SPM (the 2nd row), using the outputs from the three stages separately. Each ERF image is generated by averaging over 5000 $256\times 256$ images from ImageNet-1K validation set.
Reference Backbone #Params (M) Top-1 Acc. (%)
ICCV 2021 Swin-T Liu et al. (2021) 29.0 81.3
Swin-T (HAFA) Liu et al. (2021) 29.1 81.7
Swin-T (SPM) 29.9 82.4 (+1.1)
CVPR 2022 Shunted-T Ren et al. (2022) 11.5 79.8
Shunted-T (SPM) 11.6 80.6 (+0.8)
CVPR 2023 NAT-Mini Hassani et al. (2023) 20.0 81.8
NAT-Mini (SPM) 19.9 82.2 (+0.4)
Table 4: SPM can boost backbones with different attention mechanisms via replacing their original Patch Merging blocks. “#Params” refers to the number of parameters.
5 Ablation Study
All experiments were conducted with the random seed fixed at 0, without tuning for desired outcomes.
5.1 The Effectiveness of MSA, GLE and GTG
As presented in Tab. 5, replacing GLE with a $3\times 3$ AvgPool results in a 2.4% performance decrease, but still outperforms the baseline by 2.0%, with nearly identical parameter count and computational complexity.
Although the DWConv in the GTG module introduces a small number of learnable parameters, it has been experimentally proven to be highly effective. For instance, replacing GTG with an AvgPool of the same kernel size results in a 1.5% performance drop.
Moreover, PVT-Tiny (SPM) achieves accuracy similar to PVT-Small while using fewer FLOPs and parameters, suggesting that SPM’s advantage comes from structural design rather than resource scaling.
Backbone Impl. of Patch Merging #P (M) #F (G) Top-1 Acc. (%)
PVT-Tiny SPM (MSA + GLE) 13.6 2.7 79.5
MSA $\rightarrow$ $3\times 3$ Conv. 15.1 3.3 78.8 (-0.7)
GLE $\rightarrow$ $2\times 2$ Conv. 14.0 2.2 77.2 (-2.3)
GLE $\rightarrow$ $3\times 3$ AvgPool 12.4 2.0 77.1 (-2.4)
GTG $\rightarrow$ $7\times 7$ AvgPool 13.6 2.7 78.0 (-1.5)
$2\times 2$ Conv. 12.8 1.9 75.1
PVT-Small $2\times 2$ Conv. 24.1 3.7 79.8
Table 5: Classification results on ImageNet-1K with different patch merging strategies. “#P” and “#F” are calculated by the thop package.
As shown in Fig. 6 (a), the additional 0.8 GFLOPs overhead mainly results from the linear projection in GLE, which fuses $[Guide]$ and regular tokens via self-attention. Initially, we experimented with replacing self-attention with dynamic weighting (SKNet Li et al. (2019)), which reduced FLOPs but caused a slight drop in performance. This highlights the trade-off between efficiency and accuracy, and supports our choice of maintaining attention in GLE for better feature fusion. We are also exploring more efficient guiding mechanisms for future work.
5.2 The Effectiveness of Guide Token
We conducted comparative experiments to evaluate the effectiveness of the proposed $[Guide]$ token against two mainstream methods: the class token Dosovitskiy et al. (2020) and global average pooling (GAP) Chu et al. (2021), as presented in Tab. 6. The results demonstrate that the $[Guide]$ token improves model performance by approximately 1.7% compared to these methods, without significantly increasing the number of parameters.
Backbone Method #P (M) Top-1 Acc. (%)
PVT-Tiny GAP 14.0 77.7 (-1.8)
Class token 14.0 77.9 (-1.6)
Guide token (ours) 14.0 79.5
Table 6: Comparison between Guide token and other methods.
Furthermore, an important observation from Tab. 6 and Tab. 5 is that when the local window size of self-attention and the kernel size of convolution are both set to $2\times 2$, the self-attention method achieves higher accuracy than the convolution method. This suggests that the self-attention mechanism has a superior capability in extracting high-frequency features.
5.3 The Kernel Size of GTG
To determine the optimal kernel size for GTG, we conducted a series of performance comparison experiments with different kernel sizes, as shown in Tab. 7.
Backbone Kernel Size #P (M) Top-1 Acc. (%)
PVT-Tiny $1\times 1$ 13.6 78.17
$3\times 3$ 13.6 78.86
$5\times 5$ 13.6 79.43
$\bm{7\times 7}$ 13.6 79.47
$9\times 9$ 13.7 79.46
$31\times 31$ 14.5 79.34
Table 7: Performance comparison of GTG with different kernel sizes. “#P” is calculated by the thop package.
To assess optimization stability, we experimented with kernel sizes at extremes (1 and 31) and monitored training dynamics. As shown in Fig. 6 (b), training remains stable and converges faster than the baseline (PVT-Tiny), indicating that our $[Guide]$ token design does not introduce instability. Additionally, we had initially added residual connections to $[Guide]$ tokens, but found that removing them slightly improved performance, and the training remained stable throughout.
Figure 6: (a) FLOPs distribution across modules in SPM. (b) Stable convergence even with extreme kernel sizes (1 and 31).
5.4 Gradually Applying SPM
From Tab. 8, we conclude that 1) SPM is robust to varying input sizes containing different levels of semantic content, consistently improving performance (2nd and 3rd rows), and 2) repeated application of SPM leads to a healthy linear increase in performance (2nd, 4th and 5th rows).
Backbone Stage1 Stage2 Stage3 #P (M) Top-1 Acc. (%)
PVT-Tiny 13.2 75.1
13.3 76.4 (+1.3)
13.7 76.7 (+1.6)
13.5 78.1 (+3.0)
14.0 79.5 (+4.4)
Table 8: Gradually replacing the original patch merging in PVT-Tiny with SPM.
6 Future Work
The current SPM involves many hyperparameters that need to be considered; future work will explore a more concise and computationally efficient stepwise structure to further enhance HVTs performance, and experiments demonstrate the effectiveness of the global-guide mechanism, which will be analyzed in greater depth.
7 Conclusion
In this work, we introduced the Stepwise Patch Merging, inspired by the brain’s ability to integrate global and local information for comprehensive visual understanding. The proposed SPM framework, comprising Multi-Scale Aggregation and Guided Local Enhancement modules, demonstrates significant improvements in various computer vision tasks, including classification, detection, and segmentation. Meanwhile, we showed that SPM consistently enhances the performance of different backbone models. The robustness of SPM to different input sizes and its effective generalization further underscore its versatility and potential as a powerful enhancement technique.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 62406325).